OpenStudy (anonymous):

Consider the func y=3x^2-x^3 a) What is the equation of the tangent line to this function at the point (1,2) ?

OpenStudy (accessdenied):

Do you remember how to find the equation of a line using point-slope in Algebra? It's very much like that, only we find the slope in a different way. Point-slope: \(y - y_1 = m(x - x_1)\) \(m\) is the slope of the tangent line at \((1,2)\), or the derivative at x=3

OpenStudy (accessdenied):

x=1, not x=3 *

OpenStudy (anonymous):

To determine the equation of a line, you need SLOPE and a POINT on that line... In the question, you are given the point... How do you find the slope? (Well, in calculus, the derivative determines the rate of change, slope, of a function at a given point)

OpenStudy (anonymous):

So, take the derivative of the equation.. 6x - 3x^2... Then, plug in the x value of the point to find the slope at that point

OpenStudy (anonymous):

Taking (1, 2), plug in the x-value 1 ---- 6(1) - 3(1)^2 = 3 So, the slope is 3 and a point on the line is (1, 2)... Now, find the equation of the line!

OpenStudy (anonymous):

Thanks guys....!

OpenStudy (accessdenied):

You're welcome! :)

OpenStudy (anonymous):

good luck!